OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 4x = 15

OpenStudy (anonymous):

What is x coefficient here??

OpenStudy (anonymous):

Add 4 both the sides..

OpenStudy (anonymous):

What is (x+2)^2?

OpenStudy (anonymous):

add 4 on both sides.. x2 + 4x +4 =15 + 4 x2 + 4x +4 = 19 (x+2)2 = (squareroot 19) 2

OpenStudy (anonymous):

so 4 ?

OpenStudy (anonymous):

yaaah

OpenStudy (anonymous):

thankks !

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

Completing the square ( step by step ) Step 1 : Use b/2 = 2 where b = Coefficient of x ; in this case b=4 so b/2 = 2 Step 2 : Add (b/2)^2 to both sides in this case (B/2)^2 2^2 =4 X^2 + 4x = 15 X^2 + 4x + 4 = 15 +4 = 19 ( left side a a perfect square ) ( X + 2) ) ^2 =19 From your question the response should be (b/2)^2 = 4